OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. E(-2, 2), F(5, 1) x + _ y = - _

OpenStudy (anonymous):

Find the slope. Use the point-slope form to get the equation of the line. Convert it to standard form.

OpenStudy (anonymous):

That doesn't tell me much...

OpenStudy (neo92):

gradient(slope)(m) = (y2-y1)/(x2-x1) here in your case let's take y2=5,y1=-2,x2=1,x1=2 then, m = [5-(-2)]/[1-2] = 7/-1 = -7 Consider a general point P=(x,y) on the line, Hint : gradient of a straight line is constant. Now try to work it out.. If no good come here..

OpenStudy (anonymous):

Neo, I'm trying to find the equation, not the actual answer itself.

OpenStudy (neo92):

u need da eqn of the st line passing through those given points? isnt it?

OpenStudy (neo92):

OK.. did u try y-(-2)/x-2 = -7 y+2/x-2 = -7 y+2 = -7x + 14 y = -7x + 12

OpenStudy (anonymous):

I updated the question with what I figured out so far.

OpenStudy (neo92):

sephi.. i hav alredy givn u da answer didnt u get it..?

OpenStudy (anonymous):

Oh, there we go! Thanks Neo!

OpenStudy (anonymous):

It was X + 7y = 12

OpenStudy (neo92):

:-D